# Fault Tolerant Zero Forcing

### v2

@@ 4.12 Trees
In fact, there is a correspondence between (minimum) zero forcing sets of a tree $T$ and (minimum) path covers of $T$, obtained by taking one endpoint of each path in a path cover. This yields a zero forcing set of $T$, and all zero forcing sets are obtained in this way (see, for example, the proof of Proposition 4.2 in [AIM2008]).
We denote by $\ell(T)$ the number of vertices of degree 1 in a tree $T$, and refer to $\ell(T)$ as the leaf number of $T$. For leaky forcing, it was shown in [alameda2024leaky, Proposition 3.2] that the set of leaves of a tree with an edge is the unique 1-leaky forcing set of $T$. This implies the next result, but we provide our own proof for self containment.
@@
We know from [geneson2020reconfiguration] that every tree has a zero forcing set that consists entirely of leaves. We must show that any set of all but one leaf is a zero forcing set of $T$.
@@
If $w\not\in B$, then $\mathfrak{P}$ is compatible for $w$. If there exists some $i$ such that $\deg_{\mathcal{P}_{i}}(w)=1$, and the other endpoint of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ is in $B\setminus\{v\}$. So suppose $\deg_{\mathcal{P}_{i}}(w)=2$ for all $i$. Then both endpoints of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ is in $B\setminus\{v\}$, so we can choose a forcing starting from one endpoint of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ and terminating at $w$, while the other endpoint initiates a forcing process coloring the remainder of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$. Then as before, $w$ can force $v$ after all forces are carried out in TvT-v.
@@ 5 Graph Operations
For zero forcing, this was studied in [edholm2012] and [owens]. We begin by recalling the results therein.
Theorem 5.1 ([edholm2012],[owens]).
@@
It remains to show that $B^{\prime}\setminus\{v\}$ is a zero forcing set. Before any forces are performed, $v$ is the unique white neighbor of $x$, and therefore $x\rightarrow v$ is a valid first force. We then proceed with the forces in $\mathcal{F}^{\prime}_{v_{1}}$ to color the remaining vertices. Thus $B^{\prime}$ is a fault tolerant zero forcing set with $|B^{\prime}|\leq|B|+2$, so $\operatorname{Z}_{t}(G)\leq\operatorname{Z}_{t}(G-v)+2$. ∎
@@
We now apply our results to show that any vertex with a twin must be in every fault tolerant zero forcing set, which for certain graph families can significantly reduce the computational complexity of determining $\operatorname{Z}_{t}(G)$. We first recall the relevant results in zero forcing.
Theorem 5.11 ([hogben2022book]).
Let $G$ be a graph with (independent or adjacent) twins $u$ and $v$. Then either $u$ or $v$ is in every zero forcing set of $G$. Furthermore, $0\leq\operatorname{Z}(G)-\operatorname{Z}(G-v)\leq 1$.
Theorem 5.12.
Let $G$ be a graph with (independent or adjacent) twins $u$ and $v$. Then both $u$ and $v$ are in every fault tolerant zero forcing set. Furthermore, $0\leq\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-v)\leq 2$.
Proof.
Suppose there is a fault tolerant zero forcing set $B$ such that $u\in B$ and $v\not\in B$. Then by Theorem 5.11, $B\setminus\{u\}$ is not a zero forcing set, which contradicts the assumption that $B$ is a fault tolerant zero forcing set.
Note that if $G$ has an isolated vertex, then there is no fault tolerant zero forcing set for $G$ (2.8). Thus $\operatorname{Z}_{t}(G-v)$ exists only if deleting $v$ does not result in any isolated vertices.
Corollary 5.13.
If $G$ has a pair of (independent or adjacent) twins, then $\operatorname{Z}_{t}^{k}(G)$ does not exist for $k\geq 2$.
Proof.
A fault tolerant zero forcing set must contain both twins by Theorem 5.12, and since $k\geq 2$ at least two faults must be allowed. When two faults occur at the twins, there can be no zero forcing set. ∎
@@
Author Information
Asher Brown
Winston-Salem State University
506 Fieldstone Rd
Mooresville, NC 28115
abrown423@rams.wssu.edu
Mark Hunnell
Winston-Salem State University
2621 Audubon Dr
Winston-Salem, NC 27106
hunnellm@wssu.edu
Za’Kiyah Toomer-Sanders
Winston-Salem State University
2414 Quantum Ct
Winston Salem, NC 27106
ztoomersanders121@rams.wssu.edu
Sarah Weber
Winston-Salem State University
832 Lockland Ave
Winston Salem, NC 27103
sweber124@rams.wssu.edu
Financial Declarations
The work of the authors was partially supported by National Science Foundation Award no.2447261.

### v3

@@ 4.12 Trees
In fact, there is a correspondence between (minimum) zero forcing sets of a tree $T$ and (minimum) path covers of $T$, obtained by taking one endpoint of each path in a path cover. This yields a zero forcing set of $T$, and all zero forcing sets are obtained in this way (see, for example, the proof of Proposition 4.2 in [2]).
We denote by $\ell(T)$ the number of vertices of degree 1 in a tree $T$, and refer to $\ell(T)$ as the leaf number of $T$. For leaky forcing, it was shown in [3, Proposition 3.2] that the set of leaves of a tree with an edge is the unique 1-leaky forcing set of $T$. This implies the next result, but we provide our own proof for self containment.
@@
We know from [18] that every tree has a zero forcing set that consists entirely of leaves. We must show that any set of all but one leaf is a zero forcing set of $T$.
@@
If $w\not\in B$, then $\mathfrak{P}$ is compatible for $w$. If there exists some $i$ such that $\deg_{\mathcal{P}_{i}}(w)=1$, and the other endpoint of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ is in $B\setminus\{v\}$. So suppose $\deg_{\mathcal{P}_{i}}(w)=2$ for all $i$. Then both endpoints of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ are in $B\setminus\{v\}$, so we can choose a forcing starting from one endpoint of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ and terminating at $w$, while the other endpoint initiates a forcing process coloring the remainder of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$. Then as before, $w$ can force $v$ after all forces are carried out in TvT-v.
@@ 5 Graph Operations
For zero forcing, this was studied in [14] and [24]. We begin by recalling the results therein.
Theorem 5.1 ([14],[24]).
@@
It remains to show that $B^{\prime}\setminus\{v\}$ is a zero forcing set. Before any forces are performed, $v$ is the unique white neighbor of $x$, and therefore $x\rightarrow v$ is a valid first force. We then proceed with the forces in $\mathcal{F}^{\prime}_{v_{1}}$ to color the remaining vertices. Thus $B^{\prime}$ is a fault tolerant zero forcing set with $|B^{\prime}|\leq|B|+2$, so $\operatorname{Z}_{t}(G)\leq\operatorname{Z}_{t}(G-v)+2$.
For the other inequality, let $B$ be a minimum fault tolerant zero forcing set of $G$. Since $v$ can only force its neighbors, $B\cup N_{G}(v)$ is a fault tolerant zero forcing set of GvG-v. ∎
@@
Theorem 5.9.
Let $G$ be a graph with twins $u$ and $v$. Then $0\leq\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-v)\leq 2$.
Proof.
Note that by applying Theorem 2.11, $\operatorname{Z}_{t}(G-v)$ exists only if deleting $v$ does not result in any isolated vertices.